Find the equation (in slope-intercept form) of the line with the given slope that passes through the point with the given slope that passes through the point
with the given coordinates.

slope; 1, ordered pair. (-1,-2)

Found 2 solutions by rfer, Earlsdon:
Answer by rfer(16322)

y=mx+b
y=x-1

Answer by Earlsdon(6294)

Starting with the general form of a linear equation in the slope-intercept form:
y = mx+b Substitute m = 1 so you have:
y = (1)x+b Now to find the value of b, substitute the x- and y-coordinates fom the given point (-1,-2) into the equation above and solve for b.
-2 = -1+b Add 1 to both sides.
-1 = b so that the final equation is:
y = x-1